A group of 40 students in a University has been asked to measure their heights. The data collected are as follows:


 


Height (cm) f


135 - 139 5


140 - 144 4


145 - 149 2


150 - 154 12


155 - 159 15


160 - 164 2


 


1. What is the class' mean height?


2. Determine the mode.


3. What is the median height of the students?



1
Expert's answer
2021-11-24T15:48:25-0500

1.

x is midpoint of class, f is frequency of class

then:

mean = "\sum x_i f_i/n=\frac{137\cdot5+142\cdot4+147\cdot2+152\cdot12+157\cdot15+162\cdot2}{5+4+2+12+15+2}=151.25"


2.

mode = "L+\frac{f_m-f_{m-1}}{( f_m-f_{m-1})+(f_m-f_{m+1})}w"


where

  • L is the lower class boundary of the modal group
  • fm-1 is the frequency of the group before the modal group
  • fm is the frequency of the modal group
  • fm+1 is the frequency of the group after the modal group
  • w is the group width

modal group is 155-159


mode ="155+\frac{15-12}{( 15-12)+(15-2)}\cdot4=155.75"


3.

median = "L+\frac{n/2-B}{G}w"


where:

  • L is the lower class boundary of the group containing the median
  • n is the total number of values
  • B is the cumulative frequency of the groups before the median group
  • G is the frequency of the median group
  • w is the group width

median group is 145-149


median = "145+\frac{20-9}{2}\cdot4=123"